OpenStudy (anonymous):

Write the point slope form of the line satisfying the conditions. Then use the point slope form of the equation to write slope intercept form of the equation. In function notation slope= - 7/9, passing through (0,4)

OpenStudy (texaschic101):

point slope form : y - y1 = m(x - x1) slope(m) = -7/9 (0,4) x1 = 0 and y1 = 4 now we sub y - 4 = -7/9(x - 0) <=== point slope form y - 4 = -7/9(x - 0) -- distribute y - 4 = -7/9x -- add 4 to both sides y = -7/9x + 4 <=== slope intercept form (y = mx + b) function notation : f(x) = -7/9x + 4

OpenStudy (anonymous):

Thanks

OpenStudy (texaschic101):

anytime :)